OpenStudy (anonymous):

determine the type of number that the solution x^2+2x+5=0

OpenStudy (mathstudent55):

b^2 - 4ac = 0, 1 real solution b^2 - 4ac > 0, two real solutions b^2 - 4ac < 0, two complex solutions

OpenStudy (anonymous):

the type of number would be 1 real solution?

OpenStudy (mathstudent55):

b^2 - 4ac, a = 1, b = 2, c = 5 b^2 - 4ac = 2^2 - 4(1)(5) = 4 - 20 = -16

OpenStudy (anonymous):

so what type of number is it. this may be a ridiculous question but im very confused

OpenStudy (mathstudent55):

Since you see that b^2 - 4ac is a negative number, that means the solutions to this equation are two complex (imaginary) numbers.

OpenStudy (anonymous):

ok i get it now thanks:)